# Acceleration of Lagrangian Particles in Shell Models of Turbulence

### v1

@@ Abstract
Lagrangian acceleration has been investigated both experimentally and numerically in the past, and it has been shown to exhibit extreme fluctuations, which have been rationalized as events in which tracer particles get trapped into vortical structures such as vortex tubes or filaments. Here, we consider the statistics of acceleration within the multifractal framework, as in Biferale et al. Phys. Rev. Lett. 93 064502 (2004), and investigate the statistics of Lagrangian acceleration using shell models of turbulence, as in G. Boffetta et al. Phys. Rev. E 66, 066307 (2002), that – by construction – do not contain vortical structures. Our analysis shows that, even in the absence of coherent vortex structures, the multifractal model accurately captures the extreme intermittent fluctuations observed in the acceleration, with predictions that remain robust across a wide range of Reynolds numbers.
@@
In the Lagrangian dynamics, strong intermittency is observed in the statistical properties of the acceleration, $a_{L}$, experienced by tracer particles [16, 17, 18]. Numerical simulations and laboratory experiments have shown that extreme intermittent events of $a_{L}$ occur whenever the particle experiences the effect of vortex filaments or vortex tubes [16, 17, 19]. Actually, the dynamical time scale, $\tau_{L}$, of the Lagrangian acceleration is close to the Kolmogorov time scale, $\tau_{K}$, which characterizes the dynamics of the dissipative scales. Thus, in principle, Lagrangian acceleration is not an inertial range quantity. However, within the multifractal theory, the Kolmogorov scale $\eta_{K}$ and time $\tau_{K}$ (and thus $\tau_{L}$) are fluctuating quantities whose statistical properties are controlled by $D(h)$ [20, 21, 22], and thus inherited from the inertial range physics. It turns out that the multifractal theory provides a prediction for the probability distribution $P(a_{L})$, which depends on the Reynolds number $Re$ and the function $D(h)$ without any reference to the peculiar coherent structures of the turbulent flows [18].
Then, a simple but nontrivial question is whether or not extreme intermittent fluctuations of Lagrangian acceleration $a_{L}$ can be observed without the effect of vortex filaments or any other coherent structure in the flow, provided the presence of anomalous scaling of the structure functions. This Letter aims to answer such a question. Clearly, this cannot be addressed using the Navier-Stokes equations since vortex filaments are always present in HIT. Thus, here, we consider shell model of turbulence for modeling also single-particle Lagrangian dynamics [14], where by construction no well-defined structures such as vortex filaments do occur.
@@
the inset of Fig. 1(a) and Fig. 2(b) show an example of the time evolution of $v$ and $a$, respectively, obtained with above summation. Remarkably, both the Eulerian and Lagrangian velocity statistics of shell models are well captured by the multifractal model [14], whose basic ideas are summarized in the following.
@@
In Fig. 1(a), we show the kurtosis, $\langle|\delta_{\tau}v|^{4}\rangle/\langle|\delta_{\tau}v|^{2}\rangle^{2}$, as a function of the time lag $\tau$ for different values of the Reynolds number (viz. of the viscosity), demonstrating the strong deviations from Gaussianity due to intermittency and how they increase at decreasing the viscosity. In Fig. 1(b), using ideas from extended self-similarity [31], we show the local slopes $\hat{\xi}_{p}(\tau)=d\ln(\langle|\delta_{\tau}v|^{4}\rangle)/d\ln(\langle|\delta_{\tau}v|^{2}\rangle)$ vs $\tau$. Similarly to DNS and experimental data [15], around the Kolmogorov time, we observe a dip, which is usually interpreted as the effect of vortices (see, e.g., [32]). However, in this case, a well-defined dip is observed even if vortices are absent in the shell model. Moreover, the scaling behavior predicted by the multifractal model (horizontal line) is slowly recovered. The presence of the dip around $\tau_{K}$ was captured in [15] by using the multifractal model with a supplemental model for the crossover between dissipative and inertial time scale [33]. This crossover is model-dependent (i.e., it depends on whether we are working with the Navier-Stokes equations or the shell model), and we did not attempt to reproduce it for the shell model. However, the above observations are consistent with the fact that the dip originates in the multifractal character of the fluctuations. This observation was indeed first made in [14], where it was shown that the long crossover disappears when considering a non-intermittent version of the shell model (see Fig. 2 and Fig. 5 and the corresponding discussion in Ref. [14]). It is also worthwhile to remark that machine learning trained on Lagrangian velocity as defined in (2) has been successfully applied in predicting and classifying Lagrangian turbulence obtained by DNS data [34].
@@
Conclusions. We have shown that shell models of turbulence, despite lacking coherent vortical structures, reproduce the extreme intermittent fluctuations of Lagrangian acceleration typically observed in real turbulent flows. Using a multifractal framework, we demonstrated that the statistics of acceleration — including its moments and full probability distribution — can be accurately predicted across a wide range of Reynolds numbers using only the $D(h)$ extracted from the anomalous scaling exponents $\xi(p)$ without other fitting parameters. These results suggest that the observed intermittency in Lagrangian acceleration stems from inertial-range dynamics rather than specific geometric features such as vortex filaments. Our findings support the universality of multifractal statistics in turbulence and highlight the power of shell models as minimal yet effective tools for capturing key features also of Lagrangian turbulence.

### v2

@@ Abstract
Lagrangian acceleration has been investigated both experimentally and numerically in the past, and it has been shown to exhibit extreme fluctuations, which have been rationalized as events in which tracer particles get trapped into vortical structures such as vortex tubes or filaments. Here, we consider the statistics of acceleration within the multifractal framework, as in Biferale et al. Phys. Rev. Lett. 93 064502 (2004), and investigate the statistics of Lagrangian acceleration using shell models of turbulence, as in G. Boffetta et al. Phys. Rev. E 66, 066307 (2002), that – by construction – do not contain vortical structures. Our analysis reveals that, despite not accounting for specific coherent vortex structures, the multifractal model accurately captures the extreme intermittent fluctuations observed in the acceleration, with predictions that remain robust across a wide range of Reynolds numbers.
@@
In the Lagrangian dynamics, strong intermittency is observed in the statistical properties of the acceleration, $a_{L}$, experienced by tracer particles [16, 17, 18, 19]. Numerical simulations and laboratory experiments have shown that extreme intermittent events of $a_{L}$ occur whenever the particle experiences the effect of vortex filaments or vortex tubes [16, 17, 20]. Actually, the dynamical time scale, $\tau_{L}$, of the Lagrangian acceleration is close to the Kolmogorov time scale, $\tau_{K}$, which characterizes the dynamics of the dissipative scales. Thus, in principle, Lagrangian acceleration is not an inertial range quantity. However, within the multifractal theory, the Kolmogorov scale $\eta_{K}$ and time $\tau_{K}$ (and thus $\tau_{L}$) are fluctuating quantities whose statistical properties are controlled by $D(h)$ [21, 22, 23], and thus inherited from the inertial range physics. It turns out that the multifractal theory provides a prediction for the probability distribution $P(a_{L})$, which depends on the Reynolds number $Re$ and the function $D(h)$ without any reference to the peculiar coherent structures of the turbulent flows [18].
Then, a simple yet nontrivial question is whether or not extreme intermittent fluctuations of Lagrangian acceleration $a_{L}$ can be observed without the effect of vortex filaments or any other geometrical structure in the flow, provided the presence of anomalous scaling of the structure functions in the inertial range. This Letter aims to answer such a question. Clearly, this cannot be addressed using the Navier-Stokes equations since vortex filaments are always present in HIT. Thus, here, we consider shell model of turbulence for modeling also single-particle Lagrangian dynamics [14], where by construction no well-defined structures such as vortex filaments do occur.
@@
the inset of Fig. 1(a) and Fig. 2(b) show an example of the time evolution of $v$ and $a$, respectively, obtained with above summation. Indeed, the Lagrangian velocity can be represented as the superposition of velocity fluctuations generated by turbulent eddies, which also move with the advected tracers. As originally shown in [14], one may obtain more general representations by multiplying the shell‐model variables by suitably chosen wavelet functions. However, in the shell model, this superposition is most simply realized by the summation in Eq. (2). At any rate, $v(t)$ as defined by Eq. (2) exhibits multifractal scaling in time, as first shown in [14]. Remarkably, both the Eulerian and Lagrangian velocity statistics of shell models are well captured by the multifractal model [14], whose basic ideas are summarized in the following.
@@
In Fig. 1(a), we show the kurtosis, $\langle|\delta_{\tau}v|^{4}\rangle/\langle|\delta_{\tau}v|^{2}\rangle^{2}$, as a function of the time lag $\tau$ for different values of the Reynolds number (viz. of the viscosity), demonstrating the strong deviations from Gaussianity due to intermittency and how they increase at decreasing the viscosity. In Fig. 1(b), using ideas from extended self-similarity [32], we show the local slopes $d\ln(\langle|\delta_{\tau}v|^{4}\rangle)/d\ln(\langle|\delta_{\tau}v|^{2}\rangle)$ vs $\tau$. Similarly to DNS and experimental data [15], around the Kolmogorov time, we observe a dip, which is usually interpreted as the effect of vortices (see, e.g., [33]). However, in this case, a well-defined dip is observed even if vortices are absent in the shell model. Moreover, the scaling behavior predicted by the multifractal model (horizontal line) is slowly recovered. The presence of the dip around $\tau_{K}$ was captured in [15] by using the multifractal model with a supplemental model for the crossover between dissipative and inertial time scale [34]. This crossover is model-dependent (i.e., it depends on whether we are working with the Navier-Stokes equations or the shell model), and we did not attempt to reproduce it for the shell model. However, the above observations are consistent with the fact that the dip originates in the multifractal character of the fluctuations. This observation was indeed first made in [14], where it was shown that the long crossover disappears when considering a non-intermittent version of the shell model (see Fig. 2 and Fig. 5 and the corresponding discussion in Ref. [14]). It is also worthwhile to remark that machine learning trained on Lagrangian velocity as defined in (2) has been successfully applied in predicting and classifying Lagrangian turbulence obtained by DNS data [35].
@@
Conclusions. We have shown that shell models of turbulence, despite lacking coherent vortical structures, reproduce the extreme intermittent fluctuations of Lagrangian acceleration typically observed in real turbulent flows. Using a multifractal framework, we demonstrated that the statistics of acceleration — including its moments and full probability distribution — can be accurately predicted across a wide range of Reynolds numbers using only the $D(h)$ extracted from the anomalous scaling exponents $\xi(p)$ without other fitting parameters. Building on Refs. [36, 37], our findings may be attributed to the existence of statistical coherent structures, manifest in the shell model as a particular phase organization among the velocity modes $u_{n}$. Crucially, because interactions in shell models are strictly local in scale, any such phase‐organized structures must form through the inertial‐range energy cascade itself, and thus their development is essentially independent of the details of the dissipation mechanism. Our findings support the universality of multifractal statistics in turbulence and highlight the power of shell models as minimal yet effective tools for capturing key features also of Lagrangian turbulence.